The power supplied to a typical black-and-white television set is 88 W when the set is connected to 120 V.

(a) How much electrical energy does this set consume in 1 hour?
answer in J

(b) A color television set draws about 2.3 A when connected to 120 V. How much time is required for it to consume the same energy as the black-and-white model consumes in 1 hour?
answer in min
Answer #1

a)

E=P*t =88*3600 =316,800 J

b)

Power consumed by color telecision

P=VI=120*2.3 =276 W

since

E=P*t

=>t=E/P=316,800/276

t=1148 sec=19.1 minutes
Answer #2

a)Power = v2/R

=>given Power is 88W and v=120v then R is163.63 ohms

then Energy consumed per hour is v2 xt/R = 316800 J

b) We have energy ,Voltage,Current

Vit=Energy

t= 0.3188 hours